# Mesoscopic Modeling of High-Density Carbon Nanotube Films for Memristive Device Applications

### v1

@@ 4 Conclusion
Despite the inherent variability in initial configurations and compression protocols, our descriptor-based analysis reveals consistent trends linking morphological features to electrical transport. Curvature and buckling factors, which capture local mechanical distortions, show strong positive correlations with current, suggesting that deformation enhances intertube contact and conduction pathways. In contrast, bundling suppresses the current, most likely due to a reduction in overall connectivity across the network. Global descriptors such as CNT length and network density also play significant roles, reinforcing the idea that both local geometry and global topology jointly determine charge transport. Both Pearson and Spearman correlation matrices confirm these trends, highlighting strong monotonic and, in some cases, non-linear dependencies between structural descriptors and electrical transport. Principal component analysis further consolidates these findings, with the first component defining a structural axis characterized by high curvature, strong buckling, low bundling, and high connectivity. This reduction in dimensionality highlights that efficient transport is governed by a combination of factors rather than by any single descriptor.
@@
In summary, our results demonstrate that high-density CNT films can be rationally designed to achieve specific target morphologies by tuning the amorphous carbon content and the compression protocol, with amorphous carbon incorporated as spherical domains of controllable size and density. We also demonstrate that a limited set of structural descriptors can effectively capture and explain the electrical behavior of such films. This study lays out a foundation for a more systematic exploration of CNT-based memristive architectures and highlights the potential of mesoscale modeling as a predictive tool for guiding nanostructure design.
@@ Acknowledgments
S1 Additional figures
S2 Compression protocol
@@
S4 Correlation calculations
S4.1 Calculation of Pearson Correlations
The Pearson correlation coefficients 56 were computed to quantify the linear relationships between all pairs of structural descriptors and transport properties. Each descriptor was first normalized to zero mean and unit variance. The correlation matrix was then calculated using the standard Pearson formula:
@@
S4.2 Calculation of Spearman Rank Correlations
The Spearman rank correlation coefficients 63 were computed to capture possible non-linear but monotonic relationships between descriptors and transport quantities. For each variable, the data values were first converted into ranks, that is, replaced by their position in the ordered sequence from smallest to largest. The correlation coefficient was then evaluated as the Pearson correlation between the ranked variables:
@@
S4.3 Principal Component Analysis (PCA)
@@
Principal Component Analysis 64, 65, 66, 67, 68 was used to identify the dominant modes of structural variability and their correlation with transport properties. All structural descriptors were first normalized to have zero mean and unit variance. The covariance matrix was then constructed as:

### v2

@@ 5 Conclusions
Despite the inherent variability in initial configurations and compression protocols, our descriptor-based analysis reveals consistent trends linking morphological features to electrical transport. Curvature and buckling factors, which capture local mechanical distortions, show strong positive correlations with current, suggesting that deformation improves the contact between tubes and the effective conduction pathways. In contrast, bundling suppresses the current, most likely due to a reduction in overall connectivity across the network. Global descriptors such as CNT length and network density also play significant roles, reinforcing the idea that both local geometry and global topology jointly determine charge transport. Both Pearson and Spearman correlation matrices confirm these trends, highlighting strong monotonic and, in some cases, nonlinear dependencies between structural descriptors and electrical transport. The principal component analysis further consolidates these findings, with the first component defining a structural axis characterized by high curvature, strong buckling, low bundling, and high connectivity. This reduction in dimensionality highlights that efficient transport is governed by a concerted combination of factors rather than by any single descriptor.
@@
In summary, our results demonstrate that high-density CNT films can be rationally designed to achieve specific target morphologies by tuning the amorphous carbon content and the compression protocol, with amorphous carbon incorporated as spherical domains of controllable size and density. We also demonstrate that a limited set of structural descriptors can effectively capture and explain the electrical behavior of such films. This study lays the foundation for a more systematic exploration of structure–transport relationships in CNT-based networks and composite films and highlights the potential of mesoscale modeling as a predictive tool to guide the design of complex nanostructured materials.
@@ Acknowledgments
S1 Experimental distributions of CNT lengths and diameters in RRAM films
S2 CNT film structures obtained in this work
S3 Compression protocol
@@
S5 Correlation calculations
S5.1 Calculation of Pearson Correlations
The Pearson correlation coefficients 5 were computed to quantify the linear relationships between all pairs of structural descriptors and transport properties. Each descriptor was first normalized to zero mean and unit variance. The correlation matrix was then calculated using the standard Pearson formula:
@@
S5.2 Calculation of Spearman Rank Correlations
The Spearman rank correlation coefficients 19 were computed to capture possible non-linear but monotonic relationships between descriptors and transport quantities. For each variable, the data values were first converted into ranks, that is, replaced by their position in the ordered sequence from smallest to largest. The correlation coefficient was then evaluated as the Pearson correlation between the ranked variables:
@@
S5.3 Principal Component Analysis (PCA)
@@
Principal Component Analysis 62, 1, 24, 48, 17 was used to identify the dominant modes of structural variability and their correlation with transport properties. All structural descriptors were first normalized to have zero mean and unit variance. The covariance matrix was then constructed as: